# Preferential Latent Space Models for Networks with Textual Edges

### v1

@@ 5 Simulation
In this section, we evaluate the finite sample performance of our proposed method. We also compare with some alternative solutions, and the results are collected in the supplement. Specifically, we investigate how estimation and variable selection accuracy in simulations vary with network size $n$, the number of layers $K$, edge density and the number of edges $m_{ij}$ between nodes. In particular, we simulate data from model (1) with parameters ${\bm{a}}^{\ast}$, ${\bm{W}}^{\ast}$ and ${\bm{U}}^{\ast}$. For ${\bm{a}}^{\ast}$, we generate its entries independently from Uniform$(a_{l},a_{u})$, where $a_{l}$ and $a_{u}$ are two scalars that together modulate the density of the network; for ${\bm{U}}^{\ast}$, we generate its rows ${\bm{u}}^{\ast}_{i}$’s independently from $N_{d}({\bf 0},{\bm{I}})$, which are then scaled to ensure $\|{\bm{u}}^{\ast}_{i}\|_{2}=1$ for all $i$; for ${\bm{W}}^{\ast}$, we randomly select $q_{0}$ proportion of its entries to be nonzero and set the rest to zero; values for the nonzero entries are generated independently from Uniform(0.5, 3.5). We set $d=2$, $m_{ij}=m$, $q_{0}=0.7$ and consider $n=100,200$, $K=10,20,40,80$ and $m=1,2,4,8$. Also considered are $(a_{l},a_{u})=(-3.5,-1.8),(-3,-1),(-2,-1),(-1.4,-0.9)$, corresponding to an edge density of approximately 0.04, 0.08, 0.12 and 0.16, respectively.
@@
$\frac{\|\hat{{\bm{a}}}-{\bm{a}}^{\ast}\|_{2}^{2}}{\|{\bm{a}}^{\ast}\|^{2}_{2}},\,\,\frac{\|\hat{\bm{W}}-\bm{W}^{\ast}\|_{F}^{2}}{\|\bm{W}^{\ast}\|_{F}^{2}},\,\,\min_{{\bm{R}}:\,{\bm{R}}^{\top}{\bm{R}}={\bm{R}}{\bm{R}}^{\top}={\bm{I}}_{k}}\frac{\left\|\hat{{\bm{U}}}-{\bm{U}}^{\ast}{\bm{R}}\right\|_{F}^{2}}{\left\|{\bm{U}}^{\ast}\right\|_{F}^{2}},$
where $\hat{{\bm{a}}}$, $\hat{{\bm{W}}}$ and $\hat{{\bm{U}}}$ denote the estimators from Algorithm 1. Also reported is the relative estimation error of edge probabilities, calculated as
$\frac{1}{K}\sum_{k=1}^{K}\frac{\left\|\psi(\hat{\bm{\Theta}}^{(k)})-\psi({\bm{\Theta}}^{(k)^{\ast}})\right\|_{F}^{2}}{\left\|\psi({{{\bm{\Theta}}^{(k)}}^{\ast}})\right\|_{F}^{2}},$
where ${{\bm{\Theta}}^{(k)}}^{\ast}$ is true edge probability calculated using ${\bm{a}}^{\ast}$, ${\bm{W}}^{\ast}$ and ${\bm{U}}^{\ast}$. For the variable selection accuracy in estimating ${\bm{W}}^{\ast}$, we report the true positive rate (TPR) and false positive rate (FPR). Figures 1-3 report the estimation errors of ${{\bm{a}}}^{\ast}$, ${{\bm{W}}}^{\ast}$, ${\bm{U}}^{\ast}$ and $\psi({\bm{\Theta}}^{(k)^{\ast}})$ under various simulation settings, with 95% confidence intervals, over 100 data replications.
@@
Figure 2: Mean relative errors and their corresponding $95\%$ intervals under varying edge density, while $n=200$, $K=20$ and $m=1$.
Figure 3: Mean relative errors and their corresponding $95\%$ intervals under varying $m$, while $n=200$, $K=20$ and edge density at 0.08.
It is seen from Figure 1 that estimation errors of ${\bm{a}}^{\ast}$ and ${\bm{U}}^{\ast}$ decrease with the network size $n$ and number of layers $K$, confirming the theoretical results in Theorem 1. The relative estimation error of ${\bm{W}}^{\ast}$ does not vary with $K$, as we rescale $\|\hat{\bm{W}}-\bm{W}^{\ast}\|_{F}^{2}$ by $\|\bm{W}^{\ast}\|_{F}^{2}$ in calculating the relative error and $\|\bm{W}^{\ast}\|_{F}^{2}$ scales with $K$. Additionally, Figures 2-3 show that as edge density, modulated by ${\bm{a}}^{\ast}$, and the number of edges $m$ increase, the estimation errors of ${\bm{a}}^{\ast}$, ${\bm{W}}^{\ast}$ and ${\bm{U}}^{\ast}$ decrease.
In our experiments, the edge cross-validation procedure performs well in selecting the latent position dimension $d$ and number of nonzero entries in ${\bm{W}}$. It consistently selects the correct $d$ across all simulation settings. Tables 1-2 report the TPR and FPR in estimating the nonzero entries in ${\bm{W}}^{\ast}$, with standard errors in the parentheses, over 100 data replications. It is seen that the variable selection accuracy improves with $n$, $m$ and edge density. The selection accuracy remains relatively stable across different numbers of layers $K$’s, which is expected given that the dimension of ${\bm{W}}$ increases with $K$.
Table 1: True Positive Rate (TPR) and False Positive Rate (FPR) in estimating ${\bm{W}}^{\ast}$ under varying $n$, $K$, while $m=1$ and edge density at 0.08.
@@
Table 2: True Positive Rate (TPR) and False Positive Rate (FPR) in estimating ${\bm{W}}^{\ast}$ under varying edge density (while fixing $m=1$) and $m$ (while fixing edge density at 0.08).
@@ 6.1 Data description
The Enron email corpus [^1] https://www.cs.cmu.edu/~enron/ (Klimt and Yang, 2004), one of the most extensive publicly available datasets of its kind, contains over 500,000 emails from 158 employees spanning from November 13, 1998 to June 21, 2002. Released by the Federal Energy Regulatory Commission following its investigation of Enron, this dataset provides a unique opportunity to study the communications and organizational structures within a major corporation during various stages of its financial collapse.
The study period of this dataset can be divided into three different stages, as marked by two major events. In February 2001, Enron’s stock reached its peak and then began a dramatic decline following major sell-offs from top executives. It was later found that starting February 2001, concerns about Enron’s accounting practices were increasingly discussed internally. In October 2001, the company’s financial scandal was publicly exposed and the Securities and Exchange Commission began an investigation into Enron’s accounting practices. Accordingly, we consider three stages in our analysis: the pre-decline period from November 13, 1998, to February 1, 2001; the decline and pre-bankruptcy period from February 1, 2001, to October 1, 2001; and the bankruptcy and post-bankruptcy period from October 1, 2001, to June 21, 2002.
Our analysis focused on the emails of 154 employees whose roles and departments are documented in the dataset. These employees are from different departments including Gas (ENA Gas), Legal (ENA Legal), Enron Transportation Services (ETS), Regulatory and Government Affairs and Enron Wholesale Services (EWS), among others. Data preparation included standard text preprocessing steps such as tokenization, lower-casing, removel of punctuations and stopwords, and stemming. We then employed Latent Dirichlet Allocation (LDA) to extract keywords that inform latent topics in the corpus. Table 3 shows the 45 extracted keywords. To construct the network for each stage, we merged all emails between each pair of nodes into a single document. This process created an undirected network with $n=154$ nodes and $K=45$ keywords for each stage, denoted by $\{y_{ij}^{(k)}\}_{i,j\in[154],k\in[45]}$, where $y_{ij}^{(k)}=y_{ji}^{(k)}=1$ if the merged document between users $i$ and $j$ includes keyword $k$, and $y_{ij}^{(k)}=y_{ji}^{(k)}=0$ otherwise. This approach helps to reduce the sparsity of the network and facilitates comparison with other multi-layer network analysis methods. The edge densities of the three networks constructed for Stages 1, 2, and 3 are 0.012, 0.010, and 0.009, respectively.
@@
Table 3: Keywords extracted using LDA, sorted by their frequency of occurrences in the corpus.
@@ 6.2 Alternative approaches and link prediction
We also consider three alternative approaches when analyzing the Enron email network data:
  • Separate: This method fits a separate latent space model to each layer, that is, $\bm{\Theta}^{(k)}={\bm{a}}^{(k)}\bm{1}_{n}^{T}+\bm{1}_{n}{{\bm{a}}^{(k)}}^{\top}+{\bm{U}}^{(k)}{{\bm{U}}^{(k)}}^{\top}$ for $k\in[K]$.
  • Multiness (MacDonald et al., 2022): This method includes both a common latent structure across layers and a separate latent structure for each individual layer, written as $\bm{\Theta}^{(k)}=\bm{V}{\bm{I}}_{p_{0},q_{0}}\bm{V}^{\top}+{\bm{U}}^{(k)}{\bm{I}}_{p_{k},q_{k}}{{\bm{U}}^{(k)}}^{\top}$ for $k\in[K]$, where $\bm{V}\in\mathbb{R}^{n\times d_{0}}$ represents the matrix of common latent positions, ${\bm{U}}^{(k)}\in\mathbb{R}^{n\times d_{k}}$ collects the individual latent positions for layer $k$, and ${\bm{I}}_{p,q}=\left[\begin{array}[]{cc}{\bm{I}}_{p}&0\\ 0&-{\bm{I}}_{q}\end{array}\right].$
  • FlexMn (Zhang et al., 2020b): This method considers a common latent position $\mathbf{U}$ across layers with a layer-specific scaling matrix $\mathbf{\Lambda^{(k)}}$, written as
    $\bm{\Theta}^{(k)}={\bm{a}}^{(k)}\bm{1}_{n}^{\top}+\bm{1}_{n}{{\bm{a}}^{(k)}}^{\top}+{\bm{U}}\bm{\Lambda}^{(k)}{\bm{U}}^{\top},\text{ for }k\in[K],$
    where ${\bm{a}}^{(k)}\in\mathbb{R}^{n}$, ${\bm{U}}\in\mathbb{R}^{n\times d}$, and $\bm{\Lambda}^{(k)}\in\mathbb{R}^{d\times d}$.
To ensure a fair comparison, edge cross-validation is used in selecting the latent position dimension for all above methods. We first compare the edge prediction accuracy and then examine the estimated latent positions from different methods. Comparisons of these methods in simulations are included in the Section S7 of supplement.
@@
Network link prediction is a fundamental task in network data analysis. To compare the performance of different methods, we randomly remove 20% entries from each layer and treat them as missing data. We then apply PLSM, Multiness, FlexMn, and Separate to the remaining entries and use the fitted model to predict link probabilities for the missing entries. This procedure is repeated 100 times. Figure 4 shows the average precision-recall curves from all four methods. It is seen that Separate does not perform well as it cannot borrow information across different layers; Multiness might have suffered from over-fitting as there is a large number of sparse layers in each of the three networks.
@@ 6.3 Estimation results from different methods
We apply our proposed method, PLSM, to the three networks from Stages 1-3 and use edge cross-validation to select the latent dimension $d$ and the sparsity $s$. The latent dimension $d$ was chosen as 10 for Stages 1 and 2, and 8 for Stage 3. The number of non-zero entries in ${\bm{W}}$ was selected as 6653, 6445, and 6237 for Stages 1, 2, and 3, respectively. Among the 154 employees, 43 held director-level or higher-level positions and had significantly more email correspondence. For visualization purposes, we focus on these 43 employees in Figures 5-6. The first row in Figure 5 plots the estimated latent node positions. The estimated latent position ${\bm{u}}_{i}$’s from PLSM are of unit length, placing them on a $K$-dimensional sphere. Two nodes with closer latent positions, i.e., having a smaller angle between them, are more likely to communicate. The clear clustering pattern of nodes in Stages 1 and 2 shows that before bankruptcy, executives in different departments function relatively autonomously. In Stage 3, there is a noticeable increase in cross-departmental communications. Also in Stage 3, we observe an increase in communications involving the Legal department (nodes from the Legal department move closer to others), likely due to the legal ramifications of bankruptcy proceedings.
@@
For FlexMn, edge cross-validation selected latent dimensions of 9 for Stages 1 and 2, and 8 for Stage 3. Multiness, on the other hand, selected the maximum candidate value of 20 for the latent dimension $d$ for all three stages. The estimated latent positions from FlexMn and Multiness are shown in the second and third rows of Figure 5. For Multiness, only the common latent positions are plotted. Multiness displays similar qualitative patterns (up to a rotation) as PLSM in the estimated latent positions. In contrast, FlexMn appears less effective at distinguishing individuals from different departments in Stages 1 and 2.
@@
Figure 6: Estimated preferential latent positions $[W_{1k}{\bm{u}}_{1},\ldots,W_{nk}{\bm{u}}_{n}]^{\top}$ by PLSM on keywords “legal”, “sell”, and “cash”.
@@
In addition to the unit-length latent positions ${\bm{u}}_{i}$’s, PLSM also provides preferential latent positions calculated as $W_{ik}\mathbf{u}_{i}$’s for all $i,k$. For node $i$ on layer $k$, the direction of $W_{ik}\mathbf{u}_{i}$ is the same as ${\bm{u}}_{i}$, while its length $\|W_{ik}{\bm{u}}_{i}\|_{2}=W_{ik}$ characterizes the preference or activeness of node $i$ on layer $k$. Figure 6 shows the first three dimensions of $[W_{1k}{\bm{u}}_{1},\ldots,W_{nk}{\bm{u}}_{n}]^{\top}$. For illustration, we present the preferential latent positions for three keywords: “legal”, “sell”, and “cash.” The keyword “legal” in Enron email communications typically pertains to legal advice, agreements, regulatory issues, and discussions involving legal teams. The keyword “sell” generally involves discussions on asset transactions and marketing strategies. The keyword “cash” often relates to discussions on cash flow management and financial strategies. In Stage 1, the Legal department is more active in communications involving “legal” compared to “sell” or “cash”. Richard B. Sanders, Legal VP and Assistant General Counsel, exemplifies this focus. James D. Steffe, VP of Government Affairs and part of the Regulatory and Government Affairs (RGA) department, is seen active across all three layers in Stage 1. In contrast, Richard Shapiro, VP of Regulatory, shows less engagement in cash-related conversations, likely reflecting their differing responsibilities. John Arnold from the Gas department shows a significant interest in “sell” and engages more with the Legal and EWS departments in this context. Known as the “king of natural gas”, Arnold played a crucial role in natural gas trading and earned Enron’s largest bonus for his contributions. In Stage 2, the Gas and EWS departments are much closer compared to Stage 1, particularly in discussing legal-related issues. Aside from John Arnold, who remains actively engaged in both “sell” and “cash”, other members of the Gas and ETS departments show less activity. In Stage 3, the Legal department further increases its communication with other departments, especially in discussions related to cash-related issues. From Stages 2 to 3, there is a shift in interests within the RGA department, particularly for individuals such as Richard Shapiro (VP of Regulatory Affairs) and James D. Steffes (VP of Government Affairs). Both moved closer to other departments regarding “legal” matters, likely due to their involvement in post-bankruptcy issues. Additionally, the Legal department’s activity in “cash” discussions decreased.
@@
Figure 7: Boxplots of $W_{ik}$’s for “legal” and “cash” for nodes from different departments.
Figure 7 plots individual preferences $W_{ik}$’s for “legal” and “cash” by departments. It is seen from the left plot that the Legal and RGA departments are consistently active in legal-related issues across all three stages. Meanwhile, Gas and ETS departments had low activity in legal-related issues in Stages 1-2, and showed a noticeable increase in their activity in Stage 3, likely due to a stronger focus on legal-related issues following the bankruptcy. Additionally, within the EWS department, there is an observed increase in the variability of interest levels in legal-related matters from Stage 1 to Stage 3. The right plot of Figure 7 shows a noticeable increase in interest regarding cash-related issues in Stage 3 from the Gas, ETS, and EWS departments. Following the company’s bankruptcy, communications within these departments frequently include terms such as “cash out” and “cash crunch”. In contrast, the Legal and RGA departments show significantly decreased interest in cash-related issues in Stage 3.
One attractive feature of our method is that it can offer the above insights on node preferences for different keywords, offering detailed perspectives that are not directly available from other methods.

### v2

@@ 5 Simulation
In this section, we evaluate the finite sample performance of our proposed method. We also compare with some alternative solutions, and the results are collected in the supplement. Specifically, we investigate how estimation and variable selection accuracy in simulations vary with network size $n$, the number of layers $K$, edge density and the number of edges $m_{ij}$ between nodes. We simulate data from model (1) with parameters ${a}^{\ast}$, ${W}^{\ast}$ and ${U}^{\ast}$. For ${a}^{\ast}$, we generate its entries independently from Uniform$(a_{l},a_{u})$, where $a_{l}$ and $a_{u}$ together modulate the density of the network; for ${U}^{\ast}$, we generate its rows ${u}^{\ast}_{i}$’s independently from $N_{d}(0,{I})$, which are then scaled to ensure $\|{u}^{\ast}_{i}\|_{2}=1$ for all $i$; for ${W}^{\ast}$, we randomly select $q_{0}$ proportion of its entries to be nonzero and set the rest to zero; values for the nonzero entries are generated independently from Uniform(0.5, 3.5). We set $d=2$, $m_{ij}=m$, $q_{0}=0.7$ and consider $n=100,200$, $K=10,20,40,80$ and $m=1,2,4,8$. Also considered are $(a_{l},a_{u})=(-3.5,-1.8),(-3,-1),(-2,-1),(-1.4,-0.9)$, corresponding to an edge density of approximately 0.04, 0.08, 0.12 and 0.16, respectively.
@@
$\frac{\|\hat{{a}}-{a}^{\ast}\|_{2}^{2}}{\|{a}^{\ast}\|^{2}_{2}},\,\,\frac{\|\hat{W}-W^{\ast}\|_{F}^{2}}{\|W^{\ast}\|_{F}^{2}},\,\,\min_{R:\,{R}^{\top}{R}={R}{R}^{\top}={I}_{k}}\frac{\left\|\hat{{U}}-{U}^{\ast}R\right\|_{F}^{2}}{\left\|{U}^{\ast}\right\|_{F}^{2}},\,\,\frac{1}{K}\sum_{k=1}^{K}\frac{\left\|\psi(\hat{\Theta}^{(k)})-\psi({\Theta}^{(k)}{}^{\ast})\right\|_{F}^{2}}{\left\|\psi({{{\Theta}^{(k)}}{}^{\ast}})\right\|_{F}^{2}}.$
where $\hat{{a}}$, $\hat{{W}}$ and $\hat{{U}}$ are estimators from Algorithm 1, and $\psi({\Theta}^{(k)}{}^{\ast})$ is true edge probability calculated using ${a}^{\ast}$, ${W}^{\ast}$ and ${U}^{\ast}$.
@@
Figures 2-3 report the estimation errors of ${{a}}^{\ast}$, ${{W}}^{\ast}$, ${U}^{\ast}$ and $\psi({\Theta}^{(k)}{}^{\ast})$ under various settings, with 95% confidence intervals, over 100 data replications.
Figure 3: Mean relative errors and their corresponding $95\%$ intervals under varying edge density while $n=200$, $K=20$ and $m=1$ (top panel), and under varying $m$ while $n=200$, $K=20$ and edge density at 0.08 (bottom panel).
It is seen from Figure 2 that estimation errors of ${a}^{\ast}$ and ${U}^{\ast}$ decrease with the network size $n$ and number of layers $K$, confirming the theoretical results in Theorem 1. The relative estimation error of ${W}^{\ast}$ does not vary with $K$, as we rescale $\|\hat{W}-W^{\ast}\|_{F}^{2}$ by $\|W^{\ast}\|_{F}^{2}$ in calculating the relative error and $\|W^{\ast}\|_{F}^{2}$ scales with $K$. Additionally, Figure 3 show that as edge density, modulated by ${a}^{\ast}$, and the number of edges $m$ increase, the estimation errors of ${a}^{\ast}$, ${W}^{\ast}$ and ${U}^{\ast}$ decrease.
In Tables 1 and 2, we report the true positive rate (TPR) and false Positive rate (FPR) in estimating the nonzero entries in ${W}^{\ast}$. The results show that the variable selection accuracy improves with $n$, $m$ and edge density. The selection accuracy remains relatively stable across different numbers of layers $K$’s, which is expected since the dimension of ${W}$ increases with $K$.
Table 1: True positive rate (TPR) and false positive rate (FPR) in estimating ${W}^{\ast}$ under varying $n$, $K$, while $m=1$ and edge density at 0.08.
@@
Table 2: True positive rate (TPR) and false positive rate (FPR) in estimating ${W}^{\ast}$ under varying edge density (while fixing $m=1$) and varying $m$ (while fixing edge density at 0.08), while $n=200$ and $K=20$.
@@ 6.1 Data description
The Enron email corpus (Klimt and Yang, 2004), one of the most extensive publicly available datasets of its kind, contains over 500,000 emails from 158 employees from November 13, 1998 to June 21, 2002. Released by the Federal Energy Regulatory Commission following its investigation of Enron, this dataset provides a unique opportunity to study the communications and organizations within a major corporation during various stages of a financial collapse. The study period can be divided into three stages, as marked by two major events. In February 2001, Enron’s stock reached its peak and then began a dramatic decline following major sell-offs from top executives. It was later found that starting February 2001, concerns about Enron’s accounting practices were increasingly discussed internally. In October 2001, the company’s financial scandal was publicly exposed and the Securities and Exchange Commission began an investigation into Enron’s accounting practices. Accordingly, we consider three stages in our analysis: the pre-decline period from November, 1998, to February, 2001; the decline and pre-bankruptcy period from February, 2001, to October, 2001; and the bankruptcy and post-bankruptcy period from October, 2001, to June, 2002.
@@
Table 3: Extracted keywords sorted by their frequency of occurrences in the corpus.
Our analysis focused on the emails of 154 employees whose roles and departments are documented in the dataset. These employees are from different departments including gas, legal, transportation, regulatory and government, and wholesale, among others. Data preparation included standard text preprocessing steps such as tokenization, lower-casing, removel of punctuations and stopwords, and stemming. We then employed latent Dirichlet allocation to extract keywords that inform latent topics in the corpus. The 45 extracted keywords are shown in Table 3. To construct the network for each stage, we merged all emails between each pair of nodes into a single document. This process created an undirected network with $n=154$ nodes and $K=45$ keywords for each stage, denoted by $\{y_{ij}^{(k)}\}_{i,j\in[154],k\in[45]}$, where $y_{ij}^{(k)}=y_{ji}^{(k)}=1$ if the merged document between users $i$ and $j$ includes keyword $k$, and $y_{ij}^{(k)}=y_{ji}^{(k)}=0$ otherwise. This approach helps to reduce the sparsity of the network and facilitates comparison with other multi-layer network analysis methods. The edge densities of the three networks constructed for stages 1, 2, and 3 are 0.012, 0.010, and 0.009, respectively.
@@ 6.2 Alternative approaches and link prediction
We also consider three alternative approaches when analyzing the Enron data:
  • Separate: This method fits a separate latent space model to each layer, that is, $\Theta^{(k)}={a}^{(k)}1_{n}^{T}+1_{n}{{a}^{(k)}}^{\top}+{U}^{(k)}{{U}^{(k)}}^{\top}$ for $k\in[K]$.
  • Multiness (MacDonald et al., 2022): This method includes a common latent structure across layers and a separate latent structure for each individual layer, written as $\Theta^{(k)}=V{I}_{p_{0},q_{0}}V^{\top}+{U}^{(k)}{I}_{p_{k},q_{k}}{{U}^{(k)}}^{\top}$ for $k\in[K]$, where $V\in\mathbb{R}^{n\times d_{0}}$ is the matrix of common latent positions, ${U}^{(k)}\in\mathbb{R}^{n\times d_{k}}$ collects the individual latent positions for layer $k$, and ${I}_{p,q}=\left(\begin{array}[]{cc}{I}_{p}&0\\ 0&-{I}_{q}\end{array}\right)$.
  • FlexMn (Zhang et al., 2020): This method considers layer-specific degree heterogeneity ${a}_{k}$, and a ommon latent position ${U}$ across layers with a layer-specific scaling matrix ${\Lambda^{(k)}}$, written as $\Theta^{(k)}={a}^{(k)}1_{n}^{\top}+1_{n}{{a}^{(k)}}^{\top}+{U}{\Lambda}^{(k)}{U}^{\top}$, for $k\in[K]$, where ${a}^{(k)}\in\mathbb{R}^{n}$, ${U}\in\mathbb{R}^{n\times d}$, and ${\Lambda}^{(k)}\in\mathbb{R}^{d\times d}$.
To compare the performance of above methods in link prediction, we randomly remove 20% entries from each layer and treat them as missing data. We then apply PLSM, Multiness, FlexMn, and Separate to the remaining entries and use the fitted model to predict link probabilities for the missing entries. This procedure is repeated 100 times. To ensure a fair comparison, edge cross-validation is used in selecting the latent space dimension for all methods. Figure 4 shows the average precision-recall curves from all four methods. It is seen that Separate does not perform well as it cannot borrow information across different layers; Multiness might have suffered from over-fitting as there is a large number of sparse layers in each of the three networks. Comparisons of these methods in simulations are included in the Section S7 of supplement.
@@ 6.3 Estimation results from different methods
We apply our proposed method (PLSM) to the three networks from stages 1-3 and use edge cross-validation to select the latent dimension $d$ and the sparsity $s$. Edge cross validation selected a latent dimension of 10 for stages 1 and 2, and 8 for stage 3. The number of non-zero entries in ${W}$ was selected as 6653, 6445, and 6237 for stages 1, 2, and 3, respectively. Among the 154 employees, 43 held positions at or above director-level and had significantly more email correspondence. For visualization purposes, we focus on these 43 employees in Figures 5-6. The first row in Figure 5 plots the estimated latent node positions. The estimated latent position ${u}_{i}$’s from PLSM are of unit length, placing them on a $K$-dimensional sphere. Two nodes with closer latent positions, that is, having a smaller angle between them, are more likely to communicate. The clear clustering pattern of nodes in stages 1 and 2 shows that before bankruptcy, executives in different departments function relatively autonomously. In stage 3, there is a noticeable increase in cross-departmental communications. Also in stage 3, we observe an increase in communications involving the legal department where nodes from the legal department move closer to others, likely due to the legal ramifications of bankruptcy proceedings.
@@
For FlexMn, edge cross-validation selected latent dimensions of 9 for stages 1 and 2, and 8 for stage 3. For Multiness, edge cross-validation selected 20, the maximum candidate value, for all three stages. The estimated latent positions from FlexMn and Multiness are shown in Figure 5. For Multiness, only the common latent positions are plotted. Multiness displays similar qualitative patterns (up to a rotation) as PLSM in the estimated latent positions. In contrast, FlexMn appears less effective at distinguishing individuals from different departments in stages 1 and 2.
@@
Figure 6: Estimated preferential latent positions $W_{ik}{u}_{i}$’s by PLSM on “legal”, “sell”, and “cash”.
@@
In addition to the unit-length latent positions ${u}_{i}$’s, our proposed method also provides preferential latent positions calculated as $W_{ik}{u}_{i}$’s. In particular, the direction of $W_{ik}{u}_{i}$ is the same as ${u}_{i}$, while its length $\|W_{ik}{u}_{i}\|_{2}=W_{ik}$ characterizes the preference or activeness of node $i$ on layer $k$. Figure 6 shows the first three dimensions of $[W_{1k}{u}_{1},\ldots,W_{nk}{u}_{n}]^{\top}$ on three keywords: “legal”, “sell”, and “cash.” The keyword “legal” in Enron emails typically pertains to legal advice, agreements and regulatory issues. The keyword “sell” generally involves discussions on asset transactions and marketing strategies. The keyword “cash” often relates to discussions on cash flow management and financial strategies. In stage 1, the legal department is more active in communications involving “legal” compared to “sell” or “cash”. Richard B. Sanders, legal vice president and assistant general counsel, exemplifies this focus. James D. Steffe, vice president of government and part of the regulatory and government department, is seen active across all three layers in stage 1. In contrast, Richard Shapiro, vice president of regulatory, shows less engagement in cash-related matters, likely reflecting their differing responsibilities. John Arnold from the gas department shows a significant interest in “sell” and engages more with the legal and wholesale departments in this context. Known as the “king of natural gas”, Arnold played a crucial role in natural gas trading and earned Enron’s largest bonus for his contributions. In stage 2, the gas and wholesale departments move closer compared to stage 1, particularly in discussing legal-related issues. Aside from John Arnold, who remains actively engaged in both “sell” and “cash”, other members of the gas and transportation departments show less activity. In stage 3, the legal department further increases its communication with other departments, especially on cash-related issues. From stages 2 to 3, there is a shift in interests within the regulatory and government department, particularly for individuals such as Richard Shapiro and James D. Steffes. Both moved closer to other departments regarding legal-related matters, likely due to their involvement in post-bankruptcy issues. Additionally, the Legal department’s activity in cash-related discussions decreased.
@@
Figure 7: Boxplots of $W_{ik}$’s for “legal” and “cash” for nodes from different departments. The regulatory and government department is denoted as RGA.
Figure 7 plots node-layer preferences $W_{ik}$’s for “legal” and “cash” by departments. It is seen from Figure 7 (a) that legal, regulatory and government departments are active in legal-related issues across all three stages. Meanwhile, gas and transportation departments had low activity in legal-related issues in stages 1-2, and showed a noticeable increase in their activity in stage 3. Additionally, within the wholesale department, there is an observed increase in the variability of interest levels in legal-related matters from stage 1 to stage 3. Figure 7 (b) shows a noticeable increase in interest regarding cash-related issues in stage 3 from the gas, transportation, and wholesale departments. Following the company’s bankruptcy, communications within these departments frequently include terms such as “cash out” and “cash crunch”. In contrast, the legal and regulatory and government departments show decreased interest in cash-related issues in stage 3.
One attractive feature of our method is that it provides the above insights on node-layer preferences, offering perspectives that are not directly available from other methods.